OpenStudy (anonymous):

solve and check 2x^2+3x=14

OpenStudy (nathan917):

http://www.algebrahelp.com/calculators/equation/factoring/calc.do?equation=10x^2+-+9x+%2B+2+%3D+0 Is your friend XD Simplifying 2x2 + 3x = 14 Reorder the terms: 3x + 2x2 = 14 Solving 3x + 2x2 = 14 Solving for variable 'x'. Reorder the terms: -14 + 3x + 2x2 = 14 + -14 Combine like terms: 14 + -14 = 0 -14 + 3x + 2x2 = 0 Factor a trinomial. (-7 + -2x)(2 + -1x) = 0 Subproblem 1 Set the factor '(-7 + -2x)' equal to zero and attempt to solve: Simplifying -7 + -2x = 0 Solving -7 + -2x = 0 Move all terms containing x to the left, all other terms to the right. Add '7' to each side of the equation. -7 + 7 + -2x = 0 + 7 Combine like terms: -7 + 7 = 0 0 + -2x = 0 + 7 -2x = 0 + 7 Combine like terms: 0 + 7 = 7 -2x = 7 Divide each side by '-2'. x = -3.5 Simplifying x = -3.5 Subproblem 2 Set the factor '(2 + -1x)' equal to zero and attempt to solve: Simplifying 2 + -1x = 0 Solving 2 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + -2 + -1x = 0 + -2 Combine like terms: 2 + -2 = 0 0 + -1x = 0 + -2 -1x = 0 + -2 Combine like terms: 0 + -2 = -2 -1x = -2 Divide each side by '-1'. x = 2 Simplifying x = 2 Solution x = {-3.5, 2}

OpenStudy (anonymous):

got it thanks @nathan917